The velocity v of a particle moving in the xy plane is given by

 

v = (7.0t -6.5t2 )i + 6.0j, in m/s. Here v is in m/s and t (for positive time) is in s. What is the acceleration when t = 3.0 s? i-component of acceleration?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
As we know
acceleration (a) = dv/dt
a = (7 - 6.5 * 2 t) i + 0
a = (7 - 13t) i
Now
acceleration at t = 3sec
a = (7 - 39) = -32 m/sec²